# NTK-DFL: Enhancing Decentralized Federated Learning in Heterogeneous Settings via Neural Tangent Kernel

### v1

@@ 3.3 Weight Evolution
Lastly, while memory efficiency is not the primary focus of this paper, we briefly note a technique to address potential memory constraints in NTK-DFL implementations. For scenarios involving dense networks or large datasets, we introduce Jacobian batching. This approach allows clients to process their local datasets in smaller batches, reducing memory complexity from $O(N_{i}d_{2}d)$ to $O(N_{i}d_{2}d/m)$, where $m$ is the number of batches. Clients compute and transmit Jacobians for each batch separately, evolving their weights multiple times per communication round. This complexity reduction allows clients to connect in a denser network for the same memory cost.
4 Experiments
4.1 Experimental Setup
Datasets and Model Specifications Following Yue et al. (2022), we experiment on three datasets: Fashion-MNIST (Xiao et al., 2017), FEMNIST (Caldas et al., 2019), and MNIST (Lecun et al., 1998). Each dataset contains $C=10$ output classes. For Fashion-MNIST and MNIST, data heterogeneity has been introduced in the form of non-IID partitions created by the symmetric Dirichlet distribution (Good, 1976). For each client, a vector ${\bm{q}}_{i}\sim\text{Dir}(\alpha)$ is sampled, where ${\bm{q}}_{i}\in\mathbb{R}^{C}$ is confined to the $(C-1)$-standard simplex such that $\sum_{j=1}^{C}q_{ij}=1$. This assigns a distribution over labels to each client, creating heterogeneity in the form of label-skewness. For smaller values of $\alpha$, clients possess a distribution concentrated over fewer classes. We test over a range of $\alpha$ values in order to simulate different degrees of heterogeneity. In FEMNIST, data is split into shards based on the writer of each digit, introducing heterogeneity in the form of feature-skewness. For the model, we use a two-layer multilayer perceptron with a hidden width of 100 neurons for all trials.
Baseline Methods We compare our approach to various state-of-the-art baselines in the DFL setting. These include D-PSGD (Lian et al., 2017), DFedAvg, DFedAvgM (Sun et al., 2023), and DisPFL (Dai et al., 2022). We also compare with the centralized baseline NTK-FL (Yue et al., 2022). The upper bound NTK-FL would consist of a client fraction of $1.0$ where the servers constructs an NTK from all client data each round, which is infeasible due to memory constraints. Instead, we conducted a comparison following Dai et al. (2022), equalizing the degree of the busiest node in DFL with that of the server in centralized FL. Additional details regarding baselines and NTK-FL results can be found in Appendix B.
4.2 Experimental Results
Test Accuracy & Convergence Our experiments demonstrate the superior convergence properties of NTK-DFL compared to baselines. Figure 2 illustrates the convergence trajectories of NTK-DFL and other baselines on Fashion-MNIST. We see that NTK-DFL convergence benefits are enhanced under increased heterogeneity. Under high heterogeneity with $\alpha=0.1$, NTK-DFL establishes a 3–4% accuracy lead over the best-performing baseline within just five communication rounds and maintains this advantage throughout the training process. Additionally presented are the number of communication rounds necessary for convergence to 85% test accuracy, where NTK-DFL consistently outperforms all baselines. For the $\alpha=0.1$ setting, NTK-DFL achieves convergence in 4.6 times fewer communication rounds than DFedAvg, the next best performing baseline. Figure 8 demonstrates a similar convergence advantage for NTK-DFL on both FEMNIST and non-IID MNIST datasets.
@@ 4.2 Experimental Results
Factor Analyses of NTK-DFL We evaluate NTK-DFL’s performance over various factors, including the sparsity and heterogeneity levels, and the choices of the topology and weight initialization scheme. Figure 3 illustrates the test accuracy of NTK-DFL and other baselines as functions of the sparsity and heterogeneity levels, respectively. We observe a mild increase in convergence accuracy with decreasing sparsity. NTK-DFL experiences stable convergence across heterogeneity values $\alpha$ ranging from $0.1$ to $0.5$.
@@
The left plot reveals that NTK-DFL consistently outperforms baselines by 2–3% across all sparsity levels. The right plot demonstrates NTK-DFL’s resilience to data heterogeneity—while baseline methods’ performance deteriorates with decreasing $\alpha$, NTK-DFL maintains stable performance. Figure 10 illustrates the impact of network topology on NTK-DFL convergence. The dynamic topology accelerates convergence compared to the static topology, likely due to improved information flow among clients. Figure 10 demonstrates the effect of weight initialization on NTK-DFL performance. While random per-client initialization slightly slows convergence compared to uniform initialization, NTK-DFL exhibits robustness to these initialization differences.
@@
Figure 4: Performance gains of model averaging on convergence, trained on Fashion-MNIST. Solid lines correspond to accuracy of the aggregated global model, whereas dotted lines correspond to the mean accuracy across client models. NTK-DFL’s aggregated model maintains high performance, whereas mean client accuracy declines significantly with increased heterogeneity.
Gains Due to Final Model Aggregation  Figure 4 demonstrates the dramatic effect of final model aggregation on final test accuracy. Though the individual client models decrease in accuracy as the level of heterogeneity increases, the final aggregated model remains consistent across all levels of heterogeneity (as seen in Figures 2 and 3). In the most heterogeneous setting $\alpha=0.1$ that we tested, the difference between the mean accuracy of each client and the aggregated model accuracy is nearly 10%. A similar phenomenon is observed as the client topology becomes more sparse. For the same heterogeneity setting with a sparser topology of $\kappa=2$, the difference between these accuracies is nearly 15%. Though the individual performance of local client models may suffer under extreme conditions, the inter-client variance created by such unfavorable settings is exploited by model averaging to recuperate much of that lost performance.
Figure 7 suggests that inter-model variance enhances the performance of model averaging in DFL. While extreme dissimilarity in model weights would likely result in poor performance of the averaged model, we observe that a moderate degree of variance can be beneficial. We posit that the NTK-based update steps generate a more advantageous level of variance compared to baseline approaches, contributing to improved overall performance.
Figure 5: Final model test accuracy on Fashion-MNIST vs. the number of clients averaged for a highly heterogeneous setting with $\alpha=0.1$. The histogram shows the distribution of individual client model accuracies. Three client selection criteria are tested: (red) high-to-low (proposed), (green) random, and (blue) low-to-high.
Figure 6: Distributions of individual client model accuracy vs. the communication round for Fashion-MNIST. The proposed scheme (in red) conducts per-round averaging among neighbors, whereas the ablated setup (in blue) does not. Per-round averaging significantly reduces the skewness of the distribution of model performance.
@@
Figure 7: Relationship between model variance and final test accuracy on Fashion-MNIST. Each point represents a trial with distinct hyperparameters. The plot reveals a positive correlation between model variance and accuracy, suggesting that higher variance may benefit model averaging in DFL to a certain extent. Notably, the NTK-DFL approach demonstrates both higher accuracy and greater model variance compared to other methods.
@@
Selection Algorithm  Figure 6 demonstrates the results of the selection algorithm for the final model aggregation. The selection algorithm is highly effective in the heterogeneous setting. The effect is most notable in the $\alpha=0.1$ setting, where the performance with the selection algorithm significantly outperforms a random averaging order and the lower-bound averaging order. The proposed selection criterion requires the fewest clients to be averaged to achieve the same level of accuracy in this highly heterogeneous setting. In practical deployments, this has implications for final-round averaging in a fully decentralized setting. For example, clients may connect in a denser final topology and prioritize averaging with neighbors possessing a higher validation accuracy. This approach could optimize the efficacy of the final aggregation step while maintaining the decentralized nature of the system.
@@
Per-round Averaging Ablation Study In Figure 6, we perform an ablation study in which we remove the per-round parameter averaging that is a part of the NTK-DFL process. Here, clients forego the step of averaging their weight vectors with their neighbors during each communication round. Instead, clients compute Jacobians with respect to their original weight vector and send these to each of their neighbors (see Algorithm 3 in Appendix A). A massive distribution shift can be seen in the figure, where the distribution in the ablated setting is clearly skewed into lower accuracies. In contrast, NTK-DFL with per-round averaging demonstrates a much tighter distribution around a higher mean accuracy, effectively eliminating the long tail of low-performing models. Per-round averaging in NTK-DFL serves as a stabilizing mechanism against local model drift, safeguarding clients against convergence to suboptimal solutions early in the training process. In other words, client collaboration in the form of per-round averaging with neighbors ensures that no client lags behind in convergence. This is a particularly valuable feature in decentralized federated learning scenarios where maintaining uniformity across a diverse set of clients with heterogeneous data is a major challenge (Martínez Beltrán et al., 2023).

### v2

@@ 3.2 Proposed NTK-DFL
Lastly, while communication overhead and memory efficiency are not the primary focus of this paper, we briefly note a technique to address potential memory constraints in NTK-DFL implementations. For scenarios involving dense networks or large datasets, we introduce Jacobian batching. This approach allows clients to process their local datasets in smaller batches, reducing memory complexity from $O(N_{i}d_{2}d)$ to $O(N_{i}d_{2}d/m_{1})$, where $m_{1}$ is the number of batches. We also study communication efficiency, where NTK-DFL is resilient to compression measures such as top-$k$ sparsification and random projections. This enables significant reductions in communication costs without significantly compromising convergence. A thorough discussion of overhead can be found in Appendices D and E. We also provide results on a reconstruction attack performed using client Jacobians in Appendix F.
4 Theoretical Analysis
In this section, we derive a convergence bound for NTK-DFL by analyzing the behavior of the average client weight $\bar{{\bm{w}}}^{(k)}=\frac{1}{M}\sum_{i=1}^{M}{\bm{w}}_{i}^{(k)}$, similar to Sun et al. (2021); Shi et al. (2023). Unlike DFedAvg, the NTK-DFL bound includes a key additional dependence on the number of local iterations $T$ in its main term. The bound captures the unique advantage that NTK-DFL gains by using much larger $T$ values compared to other methods, improving convergence. The analysis also captures key factors such as the relationship between the spectral gap and model convergence, impact of data heterogeneity, and NTK approximation error.
Definition 4.1.
(DFL objective). We are interested in minimizing the global loss $\mathcal{L}({\bm{w}})$ across all clients, defined as
$\mathcal{L}(\!{\bm{w}}\!)\!=\!\sum_{i=1}^{M}\!\tfrac{N_{i}}{N_{\text{total}}}\!\mathcal{L}_{i}(\!{\bm{w}}\!),\ \mathcal{L}_{i}({\bm{w}})\!=\!\tfrac{1}{N_{i}}\!\sum_{j=1}^{N_{i}}\!\ell(\!{\bm{w}};{{\bm{x}}}_{i,j},\!\mathbf{{\bm{y}}}_{i,j}\!),$ (6)
where $\ell(\cdot)$ is a sample-wise loss applied to client data.
Definition 4.2.
(The gossip/mixing matrix). [Definition 1, (Sun et al., 2021)] The matrix $\mathbf{M}=[m_{i,j}]\in[0,1]^{m\times m}$ is assumed to satisfy the following properties: (i) (Graph) If $i\neq j$ and $(i,j)\notin\mathcal{V}$, then $m_{i,j}=0$; otherwise, $m_{i,j}>0$. (ii) (Symmetry) The matrix is symmetric, i.e., $\mathbf{M}=\mathbf{M}^{\top}$. (iii) (Null space property) The null space of $\mathbf{I}-\mathbf{M}$ is spanned by the all-ones vector, $\text{null}(\mathbf{I}-\mathbf{M})=\text{span}\{\mathbbm{1}\}$. (iv) (Spectral property) The matrix satisfies $\mathbf{I}\succeq\mathbf{M}\succ-\mathbf{I}$. The eigenvalues of $\mathbf{M}$ satisfy $1=|\lambda_{1}(\mathbf{M})|>|\lambda_{2}(\mathbf{M})|\geq\dots\geq|\lambda_{m}(\mathbf{M})|$, and the spectral gap is defined as $(1-\lambda)^{2}\in(0,1]$, where $\lambda:=\max\{|\lambda_{2}(\mathbf{M})|,|\lambda_{m}(\mathbf{M})|\}$.
Assumption 4.3.
(Standard DFL assumptions). We assume that $\nabla\mathcal{L}_{i}$ is Lipschitz continuous. We bound the variance of client gradients with $\sigma_{g}^{2}$ and the client gradient norm with $B$. We note that these assumptions are relatively mild and common in the DFL literature. Additional details can be found in Appendix G.
Assumption 4.4.
(Approximation error of NTK gradient). For any client $i\in\{1,2,\dots,M\}$, the difference between the NTK gradient and the true gradient of the loss is bounded by $\delta_{\text{NTK}}$, i.e.,
$\left\|\nabla\mathcal{L}_{i}^{\text{NTK}}({\bm{w}})-\nabla\mathcal{L}_{i}({\bm{w}})\right\|^{2}\leq\delta_{\text{NTK}}^{2},$ (7)
for all ${\bm{w}}\in\mathbb{R}^{d}$. We note that an explicit formulation of $\delta_{\text{NTK}}$ is possible given assumptions of a simplified model (Huang et al., 2021; Yue et al., 2022) (see Lemma G.9 in Appendix G). In our analysis, we treat this as a constant in order to remain model-agnostic.
Theorem 4.5.
Consider the average weight over $M$ clients $\bar{{\bm{w}}}^{(k)}\coloneqq\frac{1}{M}\sum_{i=1}^{M}{\bm{w}}_{i}^{(k)}$ for round $k\in\{1,\ldots,K\}$ and suppose the previously stated assumptions hold. Assuming that the learning rate satisfies $0<\eta\leq\frac{1}{8LT}$, we have
$\min_{1\leq k\leq K}\left\|\nabla\mathcal{L}\left(\bar{{\bm{w}}}^{(k)}\right)\right\|^{2}\leq\frac{2\bigr{[}\mathcal{L}(\bar{{\bm{w}}}^{(1)})-\mathcal{L}^{*}\bigl{]}}{K\gamma(T,\eta)}\\ +\alpha(\eta,T,\sigma_{g},\delta_{\text{NTK}})+\beta(\eta,T,\sigma_{g},\delta_{\text{NTK}},\lambda),$ (8)
where the constants are defined as
$\gamma(T,\eta)$ $\coloneqq\eta T-32\eta^{2}T^{2}L(\eta TL+1),$ (9a)
$\alpha(\eta,T,\sigma_{g},\delta_{\text{NTK}})$ $\coloneqq\tfrac{1}{\gamma(T,\eta)}2\eta T\big{[}L(\eta LT+1){}$
$\ \ \ \cdot 8\eta T(\delta_{\text{NTK}}^{2}\!+\!\sigma_{g}^{2})+\delta_{\text{NTK}}B\big{]},$ (9b)
$\beta(\eta,T,\sigma_{g},\delta_{\text{NTK}},\lambda)$ $\coloneqq\tfrac{1}{\gamma(T,\eta)}512\eta^{4}T^{4}L^{3}(\eta TL+1){}$
$\ \ \ \cdot(1-\lambda)^{-2}(\delta_{\text{NTK}}^{2}\!+\!\sigma_{g}^{2}\!+\!B^{2}).$ (9c)
Corollary 4.6.
Let the learning rate satisfy $O(1/L\sqrt{KT})$. Using similar assumptions as Theorem 4.5 in Appendix G, we have the following convergence rate for NTK-DFL:
$\min_{1\leq k\leq K}\left\|\nabla\mathcal{L}\left(\bar{{\bm{w}}}^{(k)}\right)\right\|^{2}\leq O\biggl{(}\frac{\mathcal{L}(\bar{{\bm{w}}}^{(1)})-\mathcal{L}^{*}}{\sqrt{KT}}\\ +\frac{\sqrt{T}(\delta_{\text{NTK}}B+\delta_{\text{NTK}}^{2}+\sigma_{g}^{2})}{\sqrt{K}}+\frac{T(\delta_{\text{NTK}}^{2}+\sigma_{g}^{2})}{K}\\ +\frac{T(\delta_{\text{NTK}}^{2}+\sigma_{g}^{2}+B^{2})}{(1-\lambda)^{2}K}\biggr{)}+O(\delta_{\text{NTK}}).$ (10)
From the bound above, we can see that the main convergence term can be improved when we increase the number of local iterations, $T$. However, as expected, we cannot arbitrarily increase $T$ without inducing a greater error in the terms $\delta_{\text{NTK}}$, $B$, and $\sigma_{g}$. Compared to bounds in Sun et al. (2021), $O\Big{(}\frac{1}{\sqrt{K}}+\frac{\sigma_{g}^{2}}{\sqrt{K}}+\frac{\sigma_{g}^{2}+B^{2}}{{(1-\lambda)^{2}K^{3/2}}}\Big{)}$, the first term to the right of the inequality highlights improved convergence derived from the ability to select a value of $T$ that is 1 to 2 orders of magnitude larger than that of other DFL methods. Intuitively, the bound tightens with an increasing spectral gap $(1-\lambda)^{2}$, which is associated with mixing speed. We note that there is an irreducible convergence floor from the NTK approximation term $\delta_{\text{NTK}}$. Obtaining an explicit bound on this term can be done through specific assumptions about the loss function and the model itself. For instance, as detailed in Lemma G.9, assumptions about the smoothness of $f_{i}$ and a bounded per-client residual norm lead to $\delta_{\text{NTK}}=O\big{(}K^{-1}T(\sigma_{g}^{2}+B^{2})\big{)}$, thus becoming reducible.
@@ 4 Theoretical Analysis
5 Experiments
5.1 Experimental Setup
Datasets and Model Specifications Following Yue et al. (2022), we experiment on three datasets: Fashion-MNIST (Xiao et al., 2017), FEMNIST (Caldas et al., 2019), and MNIST (Lecun et al., 1998). Each dataset contains $C=10$ output classes. For Fashion-MNIST and MNIST, data heterogeneity has been introduced in the form of non-IID partitions created by the symmetric Dirichlet distribution (Good, 1976). For each client, a vector ${\bm{q}}_{i}\sim\text{Dir}(\alpha)$ is sampled, where ${\bm{q}}_{i}\in\mathbb{R}^{C}$ is confined to the $(C-1)$-standard simplex such that $\sum_{j=1}^{C}q_{ij}=1$. This assigns a probability distribution over labels to each client, creating heterogeneity in the form of label-skewness. For smaller values of $\alpha$, a client possesses a distribution concentrated in fewer classes. We test over a range of $\alpha$ values in order to simulate different degrees of heterogeneity. In FEMNIST, data is split into shards based on the writer of each digit, introducing heterogeneity in the form of feature-skewness. For the model, we use a two-layer multilayer perceptron with a hidden width of 100 neurons for all trials.
Baseline Methods We compare our approach to various state-of-the-art baselines in the DFL setting. These include D-PSGD (Lian et al., 2017), DFedAvg, DFedAvgM (Sun et al., 2021), DFedSAM (Shi et al., 2023), and DisPFL (Dai et al., 2022). We also compare with the centralized baseline NTK-FL (Yue et al., 2022). The upper bound NTK-FL would consist of a client fraction of 1.0 where the server constructs an NTK from all client data each round, which is infeasible due to memory constraints. Instead, we conducted a comparison following Dai et al. (2022), which we include in Appendix C.1. We also include details on baseline hyperparameters in Appendix C.2.
5.2 Experimental Results
Test Accuracy & Convergence Our experiments demonstrate the superior convergence properties of NTK-DFL compared to baselines. Figure 2 illustrates the convergence trajectories of NTK-DFL and other baselines on Fashion-MNIST. We see that NTK-DFL convergence benefits are enhanced under increased heterogeneity. Under high heterogeneity with $\alpha=0.1$, NTK-DFL establishes a 3–4% accuracy lead over the best-performing baseline within just five communication rounds and maintains this advantage throughout the training process. Additionally presented are the number of communication rounds necessary for convergence to 85% test accuracy, where NTK-DFL consistently outperforms all baselines. For the $\alpha=0.1$ setting, NTK-DFL achieves convergence in 4.6 times fewer communication rounds than DFedAvg, the next best performing baseline. Figure 10 in Appendix C demonstrates a similar convergence advantage for NTK-DFL on both feature-skewed FEMNIST and label-skewed MNIST datasets.
@@ 5.2 Experimental Results
Factor Analyses for NTK-DFL We evaluate NTK-DFL’s performance over various factors, including the sparsity and heterogeneity levels, the choices of the topology, and weight initialization scheme. Figure 3 illustrates the test accuracy of NTK-DFL and other baselines as functions of the sparsity and heterogeneity levels, respectively. We observe a mild increase in convergence accuracy with decreasing sparsity. NTK-DFL experiences stable convergence across heterogeneity values $\alpha$ ranging from $0.1$ to $0.5$. The left plot reveals that NTK-DFL consistently outperforms baselines by 2–3% across all sparsity levels. The right plot demonstrates NTK-DFL’s resilience to data heterogeneity—while baseline methods’ performance deteriorates with decreasing $\alpha$, NTK-DFL maintains stable performance. In Figure 9 of Appendix C, we evaluate NTK-DFL across a range of network topologies and find that it performs consistently well across different connection structures with the same sparsity level. Additionally, Figure 12 illustrates the impact of a dynamic network topology on NTK-DFL convergence. The dynamic topology accelerates convergence compared to the static topology, likely due to improved information flow among clients. Figure 12 demonstrates the effect of weight initialization on NTK-DFL performance. While random per-client initialization slightly slows convergence compared to uniform initialization, NTK-DFL exhibits robustness to these initialization differences.
@@
Figure 4: Performance gains of model averaging on convergence, trained on Fashion-MNIST. Solid lines correspond to the accuracy of the aggregated global model, whereas dotted lines correspond to the mean accuracy across client models. NTK-DFL’s aggregated model maintains high performance, whereas mean client accuracy declines significantly with increased heterogeneity.
@@
Figure 5: Relationship between inter-client model deviation and final test accuracy on Fashion-MNIST. Each point represents a trial with distinct hyperparameters. The plot reveals a positive correlation between model deviation and accuracy, suggesting that higher deviation may benefit model averaging in DFL to a certain extent. Notably, the NTK-DFL approach demonstrates both higher accuracy and greater model deviation compared to other methods.
Gains Due to Final Model Aggregation  Figure 4 demonstrates the dramatic effect of final model aggregation on final test accuracy. Though the individual client models decrease in accuracy as the level of heterogeneity increases, the final aggregated model remains consistent across all levels of heterogeneity (as seen in Figures 2 and 3). In the most heterogeneous setting $\alpha=0.1$ that we tested, the difference between the mean accuracy of each client and the aggregated model accuracy is nearly 10%, as shown in Figure 4. A similar phenomenon is observed in Figure 9 of Appendix C as the client topology becomes more sparse. For the same heterogeneity setting with a sparser topology of $\kappa=2$, the difference between these accuracies is nearly 15%. Though the individual performance of local client models may suffer under extreme conditions, the inter-client model deviation (see detailed definition in Appendix C.3) created by such unfavorable settings is exploited by model averaging to recuperate much of that lost performance.
Figure 6: Final model test accuracy on Fashion-MNIST vs. the number of clients averaged for a highly heterogeneous setting with $\alpha=0.1$. The histogram shows the distribution of individual client model accuracies. Three client selection criteria are tested: proposed high-to-low (red), random (green), and low-to-high (blue).
Figure 7: Distributions of individual client model accuracy vs. the communication round for Fashion-MNIST. The proposed scheme (red) conducts per-round averaging among neighbors, whereas the ablated setup (blue) does not. Per-round averaging reduces the skewness of the model performance distribution.
Figure 5 suggests that inter-client model deviation enhances the performance of model averaging in DFL. While extreme dissimilarity in model weights would likely result in poor performance of the averaged model, we observe that a moderate degree of deviation can be beneficial. For example, this is seen in Li et al. (2020a) with the tuning of the “proximal term” $\mu$ that regulates this degree of client diversity in model updates. We posit that the NTK-based update steps generate a more advantageous level of deviation compared to baseline approaches, contributing to improved overall performance.
@@
Selection Algorithm  Figure 7 demonstrates the results of the selection algorithm for the final model aggregation. The “best-to-worst” selection algorithm is highly effective in a highly heterogeneous setting with $\alpha=0.1$. It significantly outperforms a random averaging order and the lower-bound averaging order, which requires the fewest clients to be averaged to achieve the same level of accuracy. In practical deployments, final model aggregation has implications in a fully decentralized setting. For example, when training starts to converge, clients may need to connect in a denser topology to shortlist neighbors with higher validation accuracies for final model aggregation.
@@
Per-round Averaging Ablation Study In Figure 7, we perform an ablation study in which we remove the per-round parameter averaging that is a part of the NTK-DFL process. Here, clients forego the step of averaging their weight vectors with their neighbors during each communication round. Instead, clients compute Jacobians with respect to their original weight vector and send these to each of their neighbors (see Algorithm 3 in Appendix A). A massive distribution shift can be seen in the figure, where the distribution in the ablated setting is clearly skewed into lower accuracies. In contrast, NTK-DFL with per-round averaging demonstrates a much tighter distribution around a higher mean accuracy, effectively eliminating the long tail of low-performing models. Per-round averaging in NTK-DFL serves as a stabilizing mechanism against local model drift, safeguarding clients against convergence to suboptimal solutions early in the training process. In other words, client collaboration in the form of per-round averaging with neighbors ensures that no client lags behind in convergence. This is a particularly valuable feature in decentralized federated learning scenarios, where maintaining uniformity across a diverse set of clients with heterogeneous data is a major challenge (Martínez Beltrán et al., 2023).
Communication and Memory Overhead While overhead is not the primary focus of our work, we propose several strategies to allow for control over these factors. To mitigate memory overhead, we experiment with Jacobian batching, where clients divide their data into batches each round and perform an NTK-DFL update per batch. As shown in Figure 14 of Appendix E, test accuracy per communication round improves with increasing batch size. To address communication overhead, we explore datapoint subsampling, where clients compute Jacobians with respect to only a fraction of their data each round. Figure 14 of Appendix E illustrates the expected accuracy drop from this approach—a natural trade-off for reduced communication. Finally, we evaluate compression techniques to reduce NTK-DFL’s communication volume further. Figure 16 of Appendix E compares the convergence of NTK-DFL with different compression methods, while Figure 16 presents convergence in terms of bits communicated. We leave readers to Appendix E for detailed discussions.